OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (texaschic101):

-2x + y = 1 ----> y = 2x + 1 now sub 2x + 1 in for y in the other equation -4x + (2x + 1) = -1 -2x + 1 = -1 -2x = -1 - 1 -2x = -2 x = 1 now sub 1 in for x in either equation to find y -2(1) + y = 1 -2 + y = 1 y = 1 + 2 y = 3 solution is : (1,3) you can even check your answer ... -4x + y = -1 -4(1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct)

OpenStudy (anonymous):

Thanks :)

OpenStudy (texaschic101):

sure thing :)